5

. Check

whether the following functions satisfy the following properties:

i) Increasing i

n both inputs.

ii) Diminishing marginal products.

iii) Constant returns to scale.

iii) Complementarity between capital and labor.


a) Y=K^0.5^N0.7


b) Y= 2K+3N


c)Y = K^0.3N^0.7


d) Y= 2N+2


Use the fact that given a function of the form Y=K^αN^β marginal products are given by

MPN=βK^αN^β−1, MPK=αK^α−1N^β


1
Expert's answer
2021-09-27T09:04:32-0400

a)"Y=K^{0.5}N0.7"

(I)"\frac{\delta Y}{\delta K}=K^{-0.5}N0.7>0\space as\space K,N>0\\

\frac{\delta Y}{\delta N}=0.7K^{0.5}N^{-0.3}>0\space as\space K,N>0"

Hence it is increasing in both inputs

ii)

"MP_K=\frac{\delta Y}{\delta K}=0.5K^{-0.5}N0.7\\\frac{\delta MP_K}{\delta K}=-(0.5)(0.5)K^{-1.5}N^{0.7}<0\\




MP_N=\frac{\delta Y}{\delta N}=0.7K^{0.5}N^{-0.3}\\\frac{\delta MP_N}{\delta N}=-(0.7)(0.3)K^{0.5}N^{-1.3}<0\\
As\space 


\frac{\delta MP_K}{\delta K}<0 \space and \space \frac{\delta MP_N}{\delta N}"

It satisfies diminishing marginal products

iii)

"Y=K^0.5N0.7"

Doubling K and N we get

"Y(2K,2N)=(2K)^{0.5}(2N)0.7\\=(2)^{1.2}K^{0.5}N0.7"

As"\space N^{1.2}" =Doubling K and N more than double Y. Therefore it does not satisfy constant returns to scale.

iv)

This is a cobb-douglas production function hence there is no complementarity between capital and labor.


b)"Y=2K+3N"

(I)

"\\
\frac{\delta Y}{\delta K}=MP_K=2>0\space as\space and \\\frac{\delta Y}{\delta N}=MP_N=3>0"

Hence it is increasing in both inputs

ii)

"\frac{\delta MP_K}{\delta K}=-(0.3)(0.7)K^{-1.7}N^{0.7}<0\\





 \frac{\delta MP_N}{\delta N}=-(-0.7)(0.3)K^{0.3}N^{-1.3}<0"

It satisfies diminishing marginal products

iii)

"Y=2K+3N"

Doubling K and N we get

"Y(2K,2N)=2(2K)+3(2N)\\=2[2K+3N]=2Y"

As doubling K and N doubles Y. Therefore it satisfies constant returns to scale.

iv)

Now, slope "\frac{MP_K}{MP_N}=\frac{2}{3}=constant"

Hence both inputs are substitutes of each other.


C)"Y=K^{0.5}N^{0.7}"

(I)

"\frac{\delta Y}{\delta K}=MP_K=0.3K^{-0.7}N^{0.7}>0\\
\frac{\delta Y}{\delta N}=MP_N=0.7K^{0.3}N^{-0.3}>0"

Hence it is increasing in both inputs

ii)

"MP_K=\frac{\delta Y}{\delta K}=0.5K^{-0.5}N0.7\\\frac{\delta MP_K}{\delta K}=-(0.5)(0.5)K^{-1.5}N^{0.7}<0\\




MP_N=\frac{\delta Y}{\delta N}=0.7K^{0.5}N^{-0.3}\\\frac{\delta MP_N}{\delta N}=-(0.7)(0.3)K^{0.5}N^{-1.3}<0\\
As\space 


\frac{\delta MP_K}{\delta K}<0 \space and \space \frac{\delta MP_N}{\delta N}"

It satisfies diminishing marginal products

iii)

"Y=K^{0.3}N^{0.7}"

Doubling K and N we get

"Y(2K,2N)=(2K)^{0.3}(2N)0.7\\=2K^{0.3}N^{0.7}\\=2Y"

Doubling K and N doubles Y. Therefore it satisfies constant returns to scale.

iv)

This is a cobb-douglas production function hence there is no complementarity between capital and labor.


D)"Y=2N+2"

(I)

"\\
\frac{\delta Y}{\delta K}=MP_K=0=0\space as\space and \\\frac{\delta Y}{\delta N}=MP_N=2>0"

Hence it is increasing in both inputs

ii)

"\frac{\delta MP_K}{\delta K}=0\\





 \frac{\delta MP_N}{\delta N}=>0"

It does not satisfy diminishing marginal products

iii)

"Y=2N+2"

Doubling K and N we get

"Y(2N)=2(2N)+2"

Therefore it does not satisfy constant returns to scale.

iv)

There is no complementarity between capital and labor